OpenStudy (anonymous):

Which set of equations would be used to solve this word problem? An airplane flew 5 hours with a 50 mph head wind. The return trip with a tail wind of the same speed took 3 hours. Find the speed of the plane in still air. d = 8(r + 50) d = 5(r + 50) and d = 3(r − 50) d = 5(r − 50) and d = 3(r + 50) None of these systems can be used to solve this problem.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

These are always tricky so let's think a moment.

OpenStudy (anonymous):

Ok

Directrix (directrix):

I'm thinking about the d = r*t formula. (distance = rate times time) and how to set it up.

OpenStudy (anonymous):

The distance formula is d = rt, where d represents distance, r represents rate, and t represents time.

OpenStudy (anonymous):

Thats what i have from my notes

Directrix (directrix):

Do you think a head wind slows the plane and a tail wind speeds the plane or the other way around? @HelloGoodmorning

OpenStudy (anonymous):

I think its this way around

Directrix (directrix):

So, the head wind slows the plane, and the tail wind speeds the plane.

Directrix (directrix):

The distance both ways on the trip are the same. You agree @HelloGoodmorning

OpenStudy (anonymous):

Yes

Directrix (directrix):

Variables: d is distance r is rate of speed in still air t is time Trip out: d is distance (r - 50) is rate of speed with head wind t = 5 hrs. Trip back: d is distance (r + 50) is rate of speed with tail wind t = 3 hours @HelloGoodmorning Do you agree with these variable assignments?

OpenStudy (anonymous):

Yes I do I see the method that you're using as wel

OpenStudy (anonymous):

But does that mean we found our answer

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

No, that is not the answer. That is part of the process.

OpenStudy (anonymous):

Oh okay.

Directrix (directrix):

Trip out: d = (r - 50)*5 Trip back: d = (r + 50)*3 @HelloGoodmorning What is the question? We need to know what to do with these two equations.

Directrix (directrix):

I think I see the correct option. Do you? @HelloGoodmorning

OpenStudy (anonymous):

No not really give me a second to re look over it

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

You are looking for something equivalent to: d = (r - 50)*5 d = (r + 50)*3 @HelloGoodmorning

OpenStudy (anonymous):

Yes i do see the answer it's b @Directrix

Directrix (directrix):

No. Look again. You have to look at every number. It's a little tricky but you can do it. @HelloGoodmorning

OpenStudy (anonymous):

ok :( lets try again

Directrix (directrix):

Option B has this in it: d = 5(r + 50) and our work did not. That is why B is not the correct option.

OpenStudy (anonymous):

oh it's d

Directrix (directrix):

No.

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

not d sorry c

OpenStudy (anonymous):

looked at it too fast

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

What might have thrown you off is the the option looks like this: d = 5(r − 50) and d = 3(r + 50) Ours looked like this: d = (r - 50)*5 d = (r + 50)*3 The commutative property says that they are the same. So, Option C is the one.

OpenStudy (anonymous):

Yes it did lol thanks very much @Directrix

Directrix (directrix):

You are welcome.